Find the measure of an angle whose measure is 18 degrees less than one half of the measure of its complement.
Answer by mestrydilip(13)

Let us assume the angle to be x.
Its complement will be 90-x
The equation will be
x-18=(90-x)/2
2(x-18)=90-x Multiply both sides of the equation by 2
2x-36=90-x
2x+x=90+36
3x=126
x=42